A force given as 2F - t - 4 = 0 acts horizontally on an object from (2,0) to (5,0). Find the impulse of the force


1
Expert's answer
2021-04-22T07:28:23-0400

Answer

Impulse can be given as

"J=\int Fdt"


"J=\int^5_2(\frac{t}{2}-2) dt\\J=[\frac{t^2}{4}-2t]^5_2\\J=-0.75N-s"